1. Calculate the volume of the gas, in liters, if 1.55 mol has a pressure of 1.20 atm at a temperature of -6 ∘C. (Answer in L)

2.Calculate the absolute temperature of the gas at which 3.33×10−3 mol occupies 498 mL at 760 torr . (Answer in K)

3.Calculate the pressure, in atmospheres, if 2.65×10−3 mol occupies 415 mL at 140 ∘C. (Answer in atm)

4.Calculate the quantity of gas, in moles, if 127 L at 60.0 ∘C has a pressure of 11.15 kPa .(Answer in moles)

Thanks!
Answer #1


(1) To find the volume of the gas Use the ideal gas law: PV = nRT p= pressure n=n moles of any gas R = gas constant T = tempe

To find the temperature of the gas Use the ideal gas law : PV = nRT T-PV nR p = 760 torr latm = 760 torr n = 333x 10-mol R =

(3) To find the pressure of the gas Use the ideal gas law: p = ? n = 2.65 x 10-mol R = 0.0821Latm/mol. K T = 140+273 = 413K V